# Studies of stationary features in jets: BL Lacertae II. Trajectory reversals and superluminal speeds on sub-parsec scales

### v1

@@ 5 Superluminal speeds of quasi-stationary component
The transverse speed $\beta^{g}_{\rm s}$ of C7 depends on the wave speed $\beta^{g}_{\rm wave}$ and its twist angle $\varphi$. The fastest wave (full line) with slopes $>0\aas@@fstack{\circ}5$ moves C7 with subrelativistic speeds $\beta^{g}_{\rm s}>0.01\,c$, which appear to be superluminal with $\beta_{\rm s}>1$ in the observer frame. In the case of the slowest transverse wave (dashed line), the wave twist angle and transverse velocity of C7 must be three times larger, $\approx 1{\aas@@fstack{\circ}5}$ and $\approx 0.03\,c$, to measure apparent speeds exceeding the speed of light. We estimate the mean and dispersion of apparent speeds of C7, $\beta_{\rm s}=9.32\pm 0.96$ and $\sigma_{\beta_{\rm s}}=8.06$, for smaller observing intervals less than a month implying that the latter describe more realistic trajectories. Then possible transverse speeds of C7 and twist angles of waves lie in the ranges $(0.08-0.28)\,c$ and $4\degr-14\degr$, respectively (see Fig. 9). It should be noted that these estimates are obtained for the case when the transverse wave propagates in the sagittal plane, which causes C7 to perform a linear reversing movement along the jet axis. An example close to such a movement is the linear reversal marked in red between 2007.36-2009.15 (Fig. 5). In fact, in most cases we observe nonlinear reversals indicating that the plane of the transverse waves makes some angle with the sagittal plane and/or the transverse waves have a twist in space. Thus, our speed estimates should be taken as evidence that relativistic transverse waves are capable of producing the apparent superluminal speeds of C7.
@@
The mean apparent speed of C7 $\overline{\beta_{\rm s}}=2.5$ corresponds to a mean speed of $\overline{\beta}^{g}_{\rm s,tr}\approx 0.05$ and $\overline{\varphi}\approx 2\aas@@fstack{\circ}5$ in the host galaxy frame (Fig. 9). Given ±$\pm 3\sigma_{\overline{\beta}_{\rm s}}\approx\pm 1$ variation in mean speed, we measure ranges of $\overline{\varphi}\approx 0\aas@@fstack{\circ}5-5\aas@@fstack{\circ}5$ and $\overline{\beta}^{g}_{\rm s,tr}\approx 0.02-0.11$. The upper limits of the twist angles and speeds of $\varphi\lesssim 12\degr$ and $\beta^{g}_{\rm s,tr}\lesssim 0.25$ (the latter is still subluminal) are given at $3\sigma_{\beta_{\rm s}}\approx 6$ level. If the apparent speed is subluminal $\beta_{\rm s}<1$ then $\varphi\lesssim 1\aas@@fstack{\circ}5$ and $\beta^{g}_{\rm s,tr}\lesssim 0.03$. Note that these estimates are model-dependent and may vary with changes in the adopted values of transverse wave velocity and jet viewing angle.
@@ 6 Discussion
The jet ridge line and moving components appear downstream from C7, the median position of which is at a distance of 0.26 mas from the 15 GHz core (Cohen et al. 2014, 2015). At high resolution based on a single epoch observation at 86 GHz the jet ridge line is detected up to a projected distance of about 0.4 mas from the core (Kim et al. 2023). We believe that the component at the core separation of approximately 0.3 mas is the C7 quasi-stationary component. The reasoning behind it is that this component is the brightest feature in the 86 GHz jet beyond 0.1 mas from the core and is situated at the expected core separation taking into account frequency-dependent core shift effect (e.g., Lobanov 1998; Sokolovsky et al. 2011; Pushkarev et al. 2012). The jet at 86 GHz has a complex structure with multiple bends within 0.4 mas from the core. This suggests that the transverse waves are generated upstream of C7 and become visible at 15 GHz downstream of C7. The magnetic fields upstream and downstream of C7 have a strong toroidal component, indicating that there is a dense spiral field at sub-parsec (Kim et al. 2023) and pc-scales (Pushkarev et al. 2023), and that the magnetic field still dominates the jet dynamics at these scales. This argues in favor of an MHD wave model (Meier et al. 2001) in which C7 is the ”master” recollimation shock wave, while the moving jet components are compressions of magnetosonic waves generated beyond the recollimation shock (Cohen et al. 2014, and references therein), and the transverse patterns of the ridge line are Alfvén MHD waves excited by C7 (Cohen et al. 2015). The latter is analogous to the excitation of a wave on a whip by shaking the handle. The C7 model proposed in section 4 assumes that the transverse waves are generated upstream of C7, somewhere between the base of the jet and C7. This diverges from the shaking whip model, where C7 is the site of wave excitation.
@@
Inhomogeneities of the jet plasma can lead to excitation of transverse waves, which move with relativistic speeds down the jet, similar to hose waves excited under high pressure. In favor of this scenario is the random nature of the twisting wave direction (or reversal direction) discussed earlier. The characteristics of the jet wave, such as amplitude, velocity, and frequency of generation, depend on the characteristics of the toroidal magnetic fields and the jet plasma.
@@
We assume that during the active state of the jet, powerful transverse waves with large amplitudes of $\lesssim 0.05$ mas are generated, which we observe as irregular C7 reversal patterns. When the plasma inhomogeneities weaken, the steady state of the jet comes in the form of quasi-periodic waves with amplitude $\lesssim 0.02$ mas during 2010-2012 (Cohen et al. 2015; Arshakian et al. 2020), which we observe as quasi-periodic C7 motion on spatial scales of $\approx 0.01$ mas between 2009.56 and 2011.70 (see the yellow line in Fig. 5 and Table 2). Modeling of relativistic MHD waves is necessary to shed light on the physical conditions of the plasma and magnetic field configuration required for the transition from the stable jet state to the active jet state, and to determine the region of wave excitation.
@@ 7 Summary
  • The overall smoothed and refined C7 trajectory is a sequence of reversible trajectories of different spatial scales. We identified 22 reversals, most of which are of R-type. Reversal motion occurs on average every 1.5 years. The number of reversals with clockwise and counterclockwise motion (13 and 9, respectively) suggests that the trajectory twist is random. The distributions of radial distances and azimuthal angles of reversal turn points are not statistically different for clockwise and counterclockwise reversals. We determined that a combination of successive reversals can form a larger RC-type reversal or a quasi-periodic trajectory identified as an RP-type reversal.
  • We propose a model in which C7 is a standing recollimation shock wave and its dynamics is governed by the passage of relativistic transverse waves (Alfvén waves) through the stationary component of C7, which moves in a plane normal to the central axis of the jet, similar to the motion of a seagull on a wave. The model assumes that relativistic transverse waves are generated upstream of C7. In the active state of the jet, irregular transverse waves with random twist are excited. The jet stabilizes into quasi-periodic waves with regular twist, which is observed as a quasi-periodic motion of C7. We estimate the mean wave frequency to be about 1.5 per year, which is the upper limit. The amplitude of the transverse waves ranges approximately from 0.01 to 0.05 mas ($0.013-0.065$ pc).
@@ Acknowledgements.
The VLBA is a facility of the National Radio Astronomy Observatory, a facility of the National Science Foundation that is operated under cooperative agreement with Associated Universities, Inc. This research has made use of data from the MOJAVE database that is maintained by the MOJAVE team (Lister et al. 2018).

### v2

@@ 5 Superluminal speeds of quasi-stationary component
The transverse speed $\beta^{g}_{\rm s}$ of C7 depends on the wave speed $\beta^{g}_{\rm wave}$ and its twist angle $\varphi$. The fastest wave (full line) with slopes $>0.5\degr$ moves C7 with subrelativistic speeds $\beta^{g}_{\rm s}>0.01\,c$, which appear to be superluminal with $\beta_{\rm s}>1$ in the observer frame. In the case of the slowest transverse wave (dashed line), the wave twist angle and transverse velocity of C7 must be three times larger, $\approx 1.5\degr$ and $\approx 0.03\,c$, to measure apparent speeds exceeding the speed of light. We estimate the mean and dispersion of apparent speeds of C7, $\beta_{\rm s}=9.32\pm 0.96$ and $\sigma_{\beta_{\rm s}}=8.06$, for smaller observing intervals less than a month implying that the latter describe more realistic trajectories. Then possible transverse speeds of C7 and twist angles of waves lie in the ranges $(0.08-0.28)\,c$ and $4\degr-14\degr$, respectively (see Fig. 9). It should be noted that these estimates are obtained for the case when the transverse wave propagates in the sagittal plane, which causes C7 to perform a linear reversing movement along the jet axis. An example close to such a movement is the linear reversal marked in red between 2007.36-2009.15 (Fig. 5). In fact, in most cases we observe nonlinear reversals indicating that the plane of the transverse waves makes some angle with the sagittal plane and/or the transverse waves have a twist in space. Thus, our speed estimates should be taken as evidence that relativistic transverse waves are capable of producing the apparent superluminal speeds of C7.
@@
The mean apparent speed of C7 $\overline{\beta_{\rm s}}=2.5$ corresponds to a mean speed of $\overline{\beta}^{g}_{\rm s,tr}\approx 0.05$ and $\overline{\varphi}\approx 2.5\degr$ in the host galaxy frame (Fig. 9). Given ±$\pm 3\sigma_{\overline{\beta}_{\rm s}}\approx\pm 1$ variation in mean speed, we measure ranges of $\overline{\varphi}\approx 0.5\degr-5.5\degr$ and $\overline{\beta}^{g}_{\rm s,tr}\approx 0.02-0.11$. The upper limits of the twist angles and speeds of $\varphi\lesssim 12\degr$ and $\beta^{g}_{\rm s,tr}\lesssim 0.25$ (the latter is still subluminal) are given at $3\sigma_{\beta_{\rm s}}\approx 6$ level. If the apparent speed is subluminal $\beta_{\rm s}<1$ then $\varphi\lesssim 1.5\degr$ and $\beta^{g}_{\rm s,tr}\lesssim 0.03$. Note that these estimates are model-dependent and may vary with changes in the adopted values of transverse wave velocity and jet viewing angle.
@@ 6 Discussion
The jet ridge line and moving components appear downstream from C7, the median position of which is at a distance of 0.26 mas from the 15 GHz core (Cohen et al. 2014, 2015). At high resolution based on a single epoch observation at 86 GHz the jet ridge line is detected up to a projected distance of about 0.4 mas from the core (Kim et al. 2023). We believe that the component at the core separation of approximately 0.3 mas is the C7 quasi-stationary component. The reasoning behind it is that this component is the brightest feature in the 86 GHz jet beyond 0.1 mas from the core and is situated at the expected core separation taking into account frequency-dependent core shift effect (e.g., Lobanov 1998; Sokolovsky et al. 2011; Pushkarev et al. 2012). The jet at 86 GHz has a complex structure with multiple bends within 0.4 mas from the core. This suggests that the transverse waves are generated upstream of C7 and become visible at 15 GHz downstream of C7. The magnetic fields upstream and downstream of C7 have a strong toroidal component, indicating that there is a dense spiral field at sub-parsec (Kim et al. 2023) and pc-scales (Pushkarev et al. 2023), and that the magnetic field still dominates the jet dynamics at these scales. This argues in favor of an MHD wave model (Meier et al. 2001) in which C7 is the ”master” recollimation shock wave, while the moving jet components are compressions of slow and fast magnetosonic waves generated beyond the recollimation shock (Cohen et al. 2014, and references therein), and the transverse patterns of the ridge line are Alfvén MHD waves excited by C7 (Cohen et al. 2015). The latter is analogous to the excitation of a wave on a whip by shaking the handle. The C7 model proposed in section 4 assumes that the transverse waves are generated upstream of C7, somewhere between the base of the jet and C7. This diverges from the shaking whip model, where C7 is the site of wave excitation.
@@
Disturbance of the plasma across the jet leads to the excitation of transverse Alfvén waves (Gurnett & Bhattacharjee 2005) that propagate downstream at relativistic speed, similar to hose waves excited under high pressure. In favor of this scenario is the random nature of the twisting wave direction (or reversal direction) discussed earlier. The characteristics of the jet wave, such as amplitude, velocity, and frequency of generation, depend on the characteristics of the toroidal magnetic fields and the jet plasma.
@@
We assume that during the active state of the jet, powerful transverse waves with large amplitudes of $\lesssim 0.05$ mas are generated, which we observe as irregular C7 reversal patterns in random directions across the jet. When the plasma disturbance weaken, the steady state of the jet comes in the form of quasi-oscillatory waves with amplitude $\lesssim 0.02$ mas during 2010-2012 (Cohen et al. 2015; Arshakian et al. 2020), which we observe as quasi-oscillatory C7 motion on spatial scales of $\approx 0.01$ mas between 2009.56 and 2011.70 (see the yellow line in Fig. 5 and Table 2). We speculate that the quasi-oscillatory waves observed during the steady state of the jet reflect the wobbling of the inner jet near the accretion disk as seen in 3-dimensional MHD simulations of the black hole-accretion disk-jet system (McKinney et al. 2013). Modeling of relativistic MHD waves is necessary to shed light on the physical conditions of the plasma and magnetic field configuration required for the transition from the stable jet state to the active jet state, and to determine the region of wave excitation.
@@ 7 Summary
  • The overall smoothed and refined C7 trajectory is a sequence of reversible trajectories of different spatial scales. We identified 22 reversals, most of which are of R-type. Reversal motion occurs on average every 1.5 years. The number of reversals with clockwise and counterclockwise motion (13 and 9, respectively) suggests that the trajectory twist is random. The distributions of radial distances and azimuthal angles of reversal turn points are not statistically different for clockwise and counterclockwise reversals. We determined that a combination of successive reversals can form a larger RC-type reversal or a quasi-oscillatory trajectory identified as an RO-type reversal.
  • To describe the reversible motion of C7, we propose a model in which C7 is a standing recollimation shock wave and its dynamics is governed by the passage of relativistic transverse waves (Alfvén waves) through the stationary component of C7, which moves in a plane normal to the central axis of the jet, similar to the motion of a seagull on a wave. In the active state of the jet, irregular transverse waves with random twist are excited. The jet stabilizes into quasi-oscillatory waves with regular twist, which is observed as a quasi-oscillatory motion of C7. We estimate the mean wave frequency to be about 1.5 per year, which is the upper limit. The amplitude of the transverse waves ranges approximately from 0.01 to 0.05 mas ($0.013-0.065$ pc).
@@
  • Important implications of the C7 model are that relativistic transverse waves are generated upstream of QSC and are mostly twisted in space, and the dynamics and kinematics of the QSC are determined by the speed, amplitude, and twist angle of the transverse wave, the maximum wave amplitude is about half of the QSC scattering size, and the superluminal speeds of the QSC indicate the presence of relativistic transverse waves in the jet.
@@ Acknowledgements.
We thank Yuri Y. Kovalev, Andrei Lobanov and Eduardo Ros for valuable comments and discussions. The VLBA is a facility of the National Radio Astronomy Observatory, a facility of the National Science Foundation that is operated under cooperative agreement with Associated Universities, Inc. This research has made use of data from the MOJAVE database that is maintained by the MOJAVE team (Lister et al. 2018). The research was supported by the YSU, in the frames of the internal grant. A.B.P. is supported in the framework of the State project “Science” by the Ministry of Science and Higher Education of the Russian Federation under the contract 075-15-2024-541.